# SQLFixAgent: Towards Semantic-Accurate SQL Generation via Multi-Agent Collaboration

### v1

@@ Introduction
Table 5 presents the ablation study of the ICL-based methods in SQLFixAgent. Starting from the complete framework of CodeS-3B + SQLFixAgent, we successively removed rubber duck debugging (COT) from SQLReviewer, Failure memory reflection (Reflexion), and multiple choice selection (MCS) from SQLRefiner and QueryCrafter. The experimental results show that each component of the SQLFixAgent framework plays a crucial role in achieving high accuracy, as their removal resulted in decreased performance across all difficulty levels.
@@
Considering that the LLM APIs charge based on token numbers, and inference time is proportional to token lengths, we emphasize the token efficiency of Text-to-SQL methods. In Figure 6, we demonstrate the average token consumption and execution accuracy of our method compare to three advanced methods, STRIKE (Nan et al. 2023), MAC-SQL and DAIL-SQL, they are all powered by GPT-3.5-turbo for a fair comparison. Since our SQLFixAgent initiates repairs only when the erroneous SQL is detected, it achieves state-of-art execution accuracy with fewer API token consumption, exhibiting higher token efficiency compared to other methods.
@@ Conclusion
In this paper, We first highlight the potential semantic error in LLM-based Text-to-SQL parsing and propose SQLFixAgent, an novel multi-agent collaborative framework designed for detecting and repairing the erroneous SQL. The extensive experiments on Bird, Spider and its three robustness variants showcase the effectiveness and universality of our method. We validated our opinion through ablation experiments, where the inexperienced foundation LLMs can serve as agents to assist the SFT-LLM in generating more accurate SQL queries. Furthermore, we underline the high token efficiency of SQLFixAgent compared to other advanced methods, providing direction for developing more practical Text-to-SQL solutions in real-world scenarios.

### v2

@@ Introduction
Table 5 presents the ablation study of the ICL-based methods in SQLFixAgent. Starting from the complete framework, we successively removed rubber duck debugging (COT) from SQLReviewer, Failure memory reflection (Reflexion), and multiple choice selection (MCS) from SQLRefiner and QueryCrafter. The experimental results show that each component of the SQLFixAgent framework is essential for generating semantically consistent results, as removing any of them led to decreased performance across all difficulty levels.
@@
Considering that the LLM APIs charge based on token number, and the inference time is proportional to token lengths, we emphasize the token efficiency of Text-to-SQL system. In Figure 6, we demonstrate the average token consumption and execution accuracy of our method compare to three advanced methods, STRIKE (Nan et al. 2023), MAC-SQL and DAIL-SQL, they are all powered by GPT-3.5-turbo for a fair comparison. Since the SQLFixAgent initiates repairs only when the erroneous SQL is detected, it achieves state-of-art execution accuracy with fewer API token consumption, exhibiting higher token efficiency compared to other methods.
@@ Conclusion
In this paper, we highlight the semantic error in LLM-based Text-to-SQL parsing and propose SQLFixAgent, a consistency-enhanced multi-agent collaborative framework designed for detecting and repairing the erroneous SQL. The extensive experiments on Bird, Spider and its three robustness variants showcase the effectiveness and universality of our method. We validated our opinion through ablation experiments, where the inexperienced foundation LLMs can serve as agents to assist the fine-tuned LLM in generating more accurate SQL queries. Furthermore, we underline the high token efficiency of SQLFixAgent compared to other advanced methods, providing direction for developing more practical Text-to-SQL solutions in real-world scenarios.